# GridShow: Omni Visual Generation

### v3

@@ 3.3 Extension Capabilities
Figure 6: Zero-shot video style transfer results. Our model incorporates characteristics from different animals (fox, red panda, tiger) while maintaining motion coherence.
Video Style Transfer
We explore zero-shot style transfer capabilities. As shown in Figure 6, our model successfully transfers distinctive features from fox, red panda, and tiger while maintaining the original temporal coherence and motion patterns.
@@
We further showcase the model’s understanding of 3D structure through zero-shot appearance editing. As demonstrated in Figure 7, given a sequence of human poses, our model can generate diverse variations by controlling attributes like hair color and clothing style through simple text. This capability stems from our layout-based training strategy, which naturally captures the relationship between spatial arrangement and semantic attributes.
These extensions validate the strong generalization potential of our approach. By leveraging the grid-based architecture and rich layout representations learned during training, GRID adapts to various downstream tasks without modifications, offering a versatile foundation for video and multi-view applications. More potential applications are demonstrated in Appendix A.7. Furthermore, since our method preserves FLUX’s original architecture, it retains all image generation capabilities of the base model, as shown in Appendix A.7.4.
@@ 4 Related Work
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) have fundamentally transformed image generation by employing iterative denoising processes to synthesize high-quality outputs. Subsequent advancements (Rombach et al., 2022; Podell et al., 2023; Ramesh et al., 2022; Saharia et al., 2022) have refined this paradigm leveraging latent spaces with significantly reduced computational costs. Diffusion Transformers (DiT) (Peebles & Xie, 2023) further advanced this area by replacing the U-Net architecture with transformer-based designs. This architectural shift improved training efficiency, paving the way for more scalable and versatile generative frameworks. Building on these, flow matching (Lipman et al., 2022; Esser et al., 2024) reformulates the generation process as a straight-path trajectory between data and noise distributions. More recently, FLUX (BlackForest, 2024), has combined the strengths of DiT and flow matching to achieve efficient and high-quality image generation. These models also integrate powerful language models (Raffel et al., 2020) and joint text-image attention mechanisms. This multimodal understanding has unlocked new possibilities for instruction-following and creative applications. Beyond generating high-quality images, text-to-image models demonstrate a strong spatial understanding that can be naturally extended to temporal dimensions through layout representations, enabling diverse downstream tasks.
@@
Diffusion-based approaches have shown remarkable progress in generalized video generation tasks (Ho et al., 2022; Blattmann et al., 2023b; Zhang et al., 2023; Blattmann et al., 2023a; He et al., 2023; Zhou et al., 2022; Wang et al., 2023a; Ge et al., 2023; Wang et al., 2023c, b; Singer et al., 2022; Zhang et al., 2023; Zeng et al., 2023). Notable works like VideoLDM (Blattmann et al., 2023b), Animatediff (Guo et al., 2023), and SVD (Chai et al., 2023) advance temporal modeling through specialized architectures. In the multi-view domain, various approaches (Watson et al., 2022; Liu et al., 2023a; Shi et al., 2023b; Long et al., 2024; Shi et al., 2023a; Lu et al., 2024; Li et al., 2023; Liu et al., 2023b; Li et al., 2024; Yang et al., 2024) focus on cross-view consistency through different attention mechanisms and feature space alignments. Recent 4D generation methods (Ren et al., 2023; Liang et al., 2024; Xie et al., 2024b; Sun et al., 2024; Wu et al., 2024) further extend to joint spatial-temporal synthesis, though often facing efficiency challenges or requiring multi-step generation. While these methods achieve remarkable results, they are typically tailored to specific tasks, relying on specialized architectures for image, video, or multi-view generation. Additionally, methods like VideoPoet (Kondratyuk et al., 2023) employ complex cross-modal alignment mechanisms to bridge different generation modes. In contrast, our approach introduces layout generation, an omni framework that transforms temporal and spatial generation into layout representations. This enables seamless multi-modal generation, to address a wide range of tasks through straightforward modifications to input representations, without the need for complex cross-modal alignment mechanisms.
@@ 5 Conclusion
We present GRID, a unified layout-based framework bridging video and multi-view generation through efficient grid representation. Our two-stage training strategy enables both robust generation and precise control, while the temporal refinement mechanism enhances motion coherence. Experiments demonstrate significant computational efficiency gains while maintaining competitive performance across tasks. The framework’s strong zero-shot generalization capabilities further enable adaptation to diverse applications without task-specific training, suggesting a promising direction for efficient visual sequence generation.
Impact Statement
This paper introduces research aimed at advancing visual sequence generation through an efficient layout-based framework. However, we must emphasize the potential risks associated with this technology, particularly in facial manipulation applications (Xie et al., 2024a; Luo et al., 2024), where our method could be misused to compromise identity security. Nevertheless, recent advances in adversarial perturbation protection mechanisms (Wan et al., 2024) provide solutions to help users protect their personal data against unauthorized model fine-tuning and malicious content generation. Therefore, we call for attention to these risks and encourage the adoption of defensive techniques to ensure the protection of personal content while advancing the development of generative AI technologies.
@@ A.1 Why Flux? Zero-shot Analysis of Foundation Models
Figure 8: Zero-shot evaluation of foundation models on grid-based multi-view generation tasks before we begin to train. Using the prompt ”a * from different angles in a mxn grid layout,” First row: Dalle3, Second row: Flux, Third row: Imagen3.
@@ 2. Data Efficiency: We achieve remarkable data efficiency improvements:
A.4 Distinction from In-Context LoRA
@@
Recent work IC-LoRA (Huang et al., 2024b, a) also utilizes grid-based layouts for image generation, which might superficially appear similar to our approach. However, a careful analysis reveals fundamental differences in both theoretical foundation and technical implementation.
Different Theoretical Foundations: The core principle of IC-LoRA is to use grid layouts as a prompt engineering technique, where multiple images are arranged together to provide in-context examples for task adaptation. This is essentially an extension of in-context learning from language models to visual domain. Their grid layout serves merely as a presentation format for example-based learning.
@@
In contrast, our approach fundamentally re-conceptualizes temporal sequences into spatial layouts. Rather than using grids for example presentation, we treat them as an inherent representation of temporal information, where spatial relationships in the grid directly correspond to temporal relationships in the sequence. This enables our model to learn and generate temporal dynamics in a holistic manner.
Distinct Technical Objectives: IC-LoRA’s technical implementation focuses on task adaptation through example pairs. Their method relies on LoRA-based fine-tuning and natural language prompts to define relationships between grid elements. However, this approach has inherent limitations in handling temporal dynamics, as it treats each grid element independently without explicit modeling of their temporal relationships.
@@
Our method, on the other hand, is specifically designed for temporal sequence generation. We introduce parallel flow-matching and dedicated temporal loss functions that explicitly model motion patterns and temporal coherence. This allows our approach to capture and generate complex temporal dynamics that are beyond the capability of example-based methods like IC-LoRA.
@@
Different Application Scopes: While IC-LoRA excels at static, example-based generation tasks through prompt engineering, it struggles with temporal sequence generation due to its fundamental design limitations. Our method, specifically designed for temporal modeling, naturally handles both static and dynamic visual generation tasks while maintaining precise control over temporal dynamics.
This analysis demonstrates that despite the superficial similarity in using grid layouts, our approach represents a fundamentally different direction in visual generation. We independently developed our method to address the specific challenges of temporal sequence generation, resulting in distinct technical contributions that go beyond the capabilities of example-based frameworks like IC-LoRA.
@@
These crucial differences are evidenced by our method’s superior performance in temporal tasks and its ability to maintain consistent motion patterns across sequences - capabilities that are fundamentally beyond the scope of IC-LoRA’s example-based approach.
A.5 Inference Details
@@
For extension tasks (style transfer, restoration, and editing), we modify the omni-inference framework to process full sequences while maintaining temporal coherence. Unlike the reference-guided generation that requires partial initialization and masking, these tasks operate on complete sequences with controlled noise injection for appearance modification.
@@ A.6 Post-Processing Pipeline
For multi-view generation results, we employ a two-stage enhancement process. First, the generated sequences are processed as video frames to ensure temporal consistency. Subsequently, we apply super-resolution using Real-ESRGAN (Wang et al., ) with anime-video-v3 weights, upscaling from 256×256 resolution to 1024×1024. This enhancement pipeline significantly improves visual quality while maintaining temporal coherence.
@@
A.7 Potential Applications
@@
A.7.1 Creative Multi-view Generation
@@
A.7.2 Flexible Frame Extension
@@
A.7.3 Future Extension to Video Understanding
Our layout-based framework shows potential in transforming traditional video understanding tasks into image-domain problems. Unlike conventional autoregressive approaches (Bai et al., 2024) that process frames sequentially, our method arranges frames in a grid layout, enabling parallel processing and global temporal modeling. This approach could benefit various video understanding tasks: for video-text retrieval, the layout representation allows direct comparison between video content and text embeddings across all frames simultaneously; for video question answering, it enables the model to attend to relevant frames across the entire sequence without sequential constraints; for video tracking and other analysis tasks, it avoids error accumulation common in traditional sequential processing. While we have not conducted specific experiments in these directions, our framework’s ability to convert temporal relationships into spatial ones through layouts offers a promising alternative to conventional video understanding paradigms, potentially enabling more efficient and effective multi-modal video analysis.
A.7.4 Maintained Image Generation Ability
@@
A.8 Limitations
@@
A.9 Multyview Camera Parameters
Building upon the dataset opensourced by Diffusion4D (Liang et al., 2024), Table 4 presents camera trajectory parameters, which serve as the foundation for consistent 4D content generation and subsequent reconstruction tasks.

### v5

@@ 3.3 Extension Capabilities
Video Motion Clone
Our framework enables natural video motion cloning through image redrawing without additional training. As demonstrated in Figure 9, we transform a cat video into videos featuring a fox, red panda, and tiger, while faithfully preserving the original motion patterns, temporal dynamics, and scene aesthetics.
@@
We demonstrate our model’s potential for practical 3D appearance editing through an innovative virtual try-on application. As shown in Figure 12, given an uncolored 3D human walking sequence from multiple viewpoints, our model can dress and style the figure through simple text prompts. This enables diverse appearance variations - from adding hair to rendering outfits - while maintaining consistent 3D structure and motion.
More results and applications are shown in Appendix A.9.
@@ 4 Related Work
Diffusion models [25, 26] have fundamentally transformed image generation by employing iterative denoising processes to synthesize high-quality outputs. Subsequent advancements [27, 28, 29, 30] have refined this paradigm leveraging latent spaces with significantly reduced computational costs. Diffusion Transformers (DiT) [31] further advanced this area by replacing the U-Net architecture with transformer-based designs. This architectural shift improved training efficiency, paving the way for more scalable and versatile generative frameworks. Building on these, flow matching [32, 1] reformulates the generation process as a straight-path trajectory between data and noise distributions. More recently, FLUX [33], has combined the strengths of DiT and flow matching to achieve efficient and high-quality image generation. These models also integrate powerful language models [34] and joint text-image attention mechanisms. This multimodal understanding has unlocked new possibilities for instruction-following and creative applications. Beyond generating high-quality images, text-to-image models demonstrate a strong spatial understanding that can be naturally extended to temporal dimensions through layout representations, enabling diverse downstream tasks.
@@
Diffusion-based approaches have shown remarkable progress in generalized video generation tasks [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 37, 46, 20]. Notable works like VideoLDM [36], Animatediff [18], and SVD [47] advance temporal modeling through specialized architectures. In the multi-view domain, various approaches [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 14, 17] focus on cross-view consistency through different attention mechanisms and feature space alignments. Recent 4D generation methods [58, 13, 59, 60, 61] further extend to joint spatial-temporal synthesis, though often facing efficiency challenges or requiring multi-step generation. While these methods achieve remarkable results, they are typically tailored to specific tasks, relying on specialized architectures for image, video, or multi-view generation. Additionally, methods like VideoPoet [62] employ complex cross-modal alignment mechanisms to bridge different generation modes. In contrast, our approach introduces layout generation, an omni framework that transforms temporal and spatial generation into layout representations. This enables seamless multi-modal generation, to address a wide range of tasks through straightforward modifications to input representations, without the need for complex cross-modal alignment mechanisms.
@@ 5 Conclusion
We present GRID, an omni visual generation framework through grid representation. Our two-stage training strategy enables both robust generation and precise control, while the temporal refinement mechanism enhances motion coherence. Experiments demonstrate significant computational efficiency gains while maintaining competitive performance across tasks. The framework’s strong zero-shot generalization capabilities further enable adaptation to diverse applications without task-specific training, suggesting a promising direction for efficient visual sequence generation.
@@ A.1 Why Flux? Zero-shot Analysis of Foundation Models
Figure 4: Zero-shot evaluation of foundation models on grid-based multi-view generation tasks before we begin to train. Using the prompt ”a * from different angles in a mxn grid layout,”
@@ 2. Data Efficiency: We achieve remarkable data efficiency improvements:
A.4 Distinction from Grid-based Methods
@@
Several recent works utilize grid-based layouts for image generation, including IC-LoRA [63, 64], Instant3D [65], DSD [66], OmniControl [67]. While these approaches might superficially appear similar to ours, a careful analysis reveals fundamental differences in both theoretical foundation and technical implementation.
Different Theoretical Foundations: Prior grid-based methods primarily focus on appearance consistency in static image generation. For instance, Instant3D utilizes grids to generate consistent multi-view images but lacks dedicated modeling of temporal dynamics. Similarly, DSD leverages grid structures for self-distillation to improve image quality, while OmniControl introduces Spatial Control Signals to achieve static control but does not address temporal relationships. IC-LoRA uses grid layouts as a prompt engineering technique, where multiple images are arranged together to provide in-context examples for task adaptation.
@@
In contrast, our approach fundamentally re-conceptualizes temporal sequences into spatial layouts. Rather than using grids for static consistency or example presentation, we treat them as an inherent representation of temporal information, where spatial relationships in the grid directly correspond to temporal relationships in the sequence. This enables our model to learn and generate temporal dynamics in a holistic manner.
Distinct Technical Objectives: While Flux shows potential for grid-based generation, it exhibits several limitations. As demonstrated in our experiments (Figure 8), Flux struggles with understanding grid structures, leading to inconsistencies in structure and appearance, poor performance in cross-frame object consistency, and unreliable responses to motion instructions. Other methods like IC-LoRA rely on LoRA-based fine-tuning and natural language prompts to define relationships between grid elements, treating each grid element independently without explicit modeling of their temporal relationships.
@@
Our method, specifically designed for temporal sequence generation, introduces parallel flow-matching and dedicated temporal loss functions that explicitly model motion patterns and temporal coherence. This allows our approach to capture and generate complex temporal dynamics that are beyond the capability of existing grid-based methods. We efficiently incorporate diverse data and world knowledge (3D, video) through our omni paradigm, transcending the limitations of traditional grid methods that only focus on appearance consistency.
@@
Different Application Scopes: While prior grid-based methods excel at static image generation or limited motion capabilities, they struggle with temporal sequence generation due to fundamental design limitations. Our method naturally handles both static and dynamic visual generation tasks while maintaining precise control over temporal dynamics, supporting a wide range of applications from text-to-video and multi-view generation to video editing.
@@
These crucial differences are evidenced by our method’s superior performance in temporal tasks and its ability to maintain consistent motion patterns across sequences - capabilities that are fundamentally beyond the scope of existing grid-based approaches.
A.5 Implementation Details
Training Configurations: For training our model across different tasks, we employ the following configuration details:
  • Video Generation (High Resolution): 4×4 grid format with single frame resolution of 512×768, resulting in a total resolution of 2048×3072
  • Video Generation (Extended Frames): 8×8 grid format with single frame resolution of 256×384, resulting in a total resolution of 2048×3072
  • Multi-view Generation: 4×6 grid format with single frame resolution of 256×256, resulting in a total resolution of 1024×1536
Model Variant Details:
  • Stage1: Trained only on large-scale data like WebVid and TikTok, using automatically generated descriptions
  • Stage2: Fine-tuned on Stage1 using high-quality descriptions generated by GPT-4, but without temporal loss
  • Full: Built upon Stage2 with the addition of temporal loss ($\mathcal{L}_{flow}$)
@@ A.5 Implementation Details
Extension Inference Details: For extension tasks (style transfer, restoration, and editing), we modify the omni-inference framework to process full sequences while maintaining temporal coherence. Unlike the reference-guided generation that requires partial initialization and masking, these tasks operate on complete sequences with controlled noise injection for appearance modification.
@@ A.6 Post-Processing Pipeline
For multi-view generation results, we employ a two-stage enhancement process. First, the generated sequences are processed as video frames to ensure temporal consistency. Subsequently, we apply super-resolution using Real-ESRGAN [68] with anime-video-v3 weights, upscaling from 256×256 resolution to 1024×1024. This enhancement pipeline significantly improves visual quality while maintaining temporal coherence.
@@
A.7 Extended Video Generation Capabilities
Maximum Frame Count and Resolution: GRID demonstrates significant scalability in video generation tasks. Our framework supports up to 64 frames (2.7 seconds at 24 FPS) in a single generation pass, substantially outperforming mainstream methods like AnimateDiff V3 and VideoCrafter2 which are typically limited to 16 frames. For resolution flexibility, GRID supports single-frame resolutions of 512×768 for 16-frame generation and 256×384 for 64-frame generation.
Comparison with Existing Methods: Table 5 provides a comprehensive comparison of maximum frame count capabilities across different video generation models. While CogVideo supports up to 81 frames and Wonder 2.1 similarly supports 81 frames, our qualitative analysis reveals that frame quality and semantic consistency in these models often degrade significantly beyond 40-50 frames, particularly for complex scenes and motions.
Table 5: Maximum single-pass frame count comparison across video generation models
Model Max Frames Default Resolution
AnimateDiff V3 16 256×384
VideoCrafter2 16 256×384
GRID (Ours) 64 256×384
CogVideo 81 720×480
Wonder 2.1 81 512×512
Zero-shot Extension Capability: A unique advantage of GRID is its inherent zero-shot extension capability. As demonstrated in Figure 11 of our main paper, models trained on 4×4 grid formats can directly extend to 4×8 formats without any additional training. This enables extending videos to 128 frames (5 seconds) while maintaining both semantic consistency and visual quality.
Unlike traditional diffusion methods that suffer from semantic drift during extension, our approach demonstrates:
  • Superior temporal consistency across extended sequences
  • Better frame-to-frame coherence in action and motion
  • Maintained visual quality and semantic alignment throughout extended videos
Furthermore, this capability can be combined with our sequence splicing strategy to generate even longer videos if needed for specific applications. It’s worth noting that for all existing methods, generating more than 81 frames while maintaining high quality and semantic consistency is a significant challenge. Claims of hundred-frame capabilities in some models often rely on additional frame interpolation or splicing techniques, which should be distinguished from native single-model capabilities.
Resource Efficiency Scaling: An important aspect of GRID’s design is its exceptional resource efficiency while maintaining quality. Compared to leading models, GRID achieves:
  • 6-67× faster inference acceleration
  • Approximately 1000× reduction in training data requirements (35K videos vs. 35M videos)
  • 90-97
This efficiency advantage becomes even more pronounced when generating longer videos, making GRID particularly suitable for resource-constrained environments or real-time applications.
A.8 Additional Benchmark Evaluations
T2VCompBench Evaluation: To comprehensively evaluate our method’s capabilities in video generation, we conducted evaluations on T2VCompBench [69], which measures various dimensions of video generation quality. Table 6 presents the results.
Table 6: Evaluation results on T2VCompBench benchmark
[…]
OpenSora 1.2 0.5639 0.0189 0.5063 0.2468 0.4833 0.5039 0.3719 12.0 $>$35M
CogVideoX-5B 0.6164 0.0219 0.5172 0.2658 0.5333 0.6069 0.3706 17.8 35M
GRID (Stage2) 0.5669 0.0098 0.4711 0.1987 0.5103 0.4512 0.3579 7.2 35K
GRID (Ours) 0.6132 0.0114 0.4918 0.2175 0.5342 0.4829 0.3686 7.2 35K
Our GRID model achieves comparable or even superior performance to specialized large models in attribute consistency (0.6132) and action expression (0.5342). The significant improvement from Stage2 (0.5669) to Full model (0.6132) in Consist-attr score demonstrates the effectiveness of our temporal loss in enhancing inter-frame consistency.
Furthermore, GRID demonstrates remarkable efficiency advantages:
  • Computational efficiency: 1.7× faster inference than OpenSora and 2.5× faster than CogVideoX
  • Parameter efficiency: 91% fewer parameters than OpenSora and 97% fewer than CogVideoX
  • Data efficiency: Training data requirements are 1000× less than existing methods
Inter-frame Consistency Evaluation: To specifically measure the improvement in inter-frame consistency provided by our temporal loss function, we conducted a detailed comparison between our Stage2 and Full models. As shown in Table 7, the addition of our temporal loss ($\mathcal{L}_{flow}$) significantly improves consistency measures.
Table 7: Inter-frame consistency comparison between model variants
[…]
Ours (Stage2) 0.5669 0.7214
Ours (Full) 0.6132 0.8567
The significant improvement in both Consist-attr (from T2VCompBench) and our Flow Consistency Score demonstrates that our temporal loss effectively enhances the model’s ability to maintain consistent object attributes and motion patterns across frames, addressing a key limitation of existing methods.
A.9 Potential Applications
@@
A.9.1 Creative Multi-view Generation
@@
Figure 9: Zero-shot video motion clone results. Our model incorporates characteristics from different animals (fox, red panda, tiger) while maintaining motion pattern.
@@
A.9.2 Flexible Frame Extension
@@
A.9.3 Future Extension to Video Understanding
Our layout-based framework shows potential in transforming traditional video understanding tasks into image-domain problems. Unlike conventional autoregressive approaches [70] that process frames sequentially, our method arranges frames in a grid layout, enabling parallel processing and global temporal modeling. This approach could benefit various video understanding tasks: for video-text retrieval, the layout representation allows direct comparison between video content and text embeddings across all frames simultaneously; for video question answering, it enables the model to attend to relevant frames across the entire sequence without sequential constraints; for video tracking and other analysis tasks, it avoids error accumulation common in traditional sequential processing. While we have not conducted specific experiments in these directions, our framework’s ability to convert temporal relationships into spatial ones through layouts offers a promising alternative to conventional video understanding paradigms, potentially enabling more efficient and effective multi-modal video analysis.
A.9.4 Maintained Image Generation Ability
@@
A.10 Impact Statement
This paper introduces research aimed at advancing visual sequence generation through an efficient layout-based framework. However, we must emphasize the potential risks associated with this technology [71], particularly in facial manipulation applications [72, 73], where our method could be misused to compromise identity security. Nevertheless, recent advances in adversarial perturbation protection mechanisms [74] provide solutions to help users protect their personal data against unauthorized model fine-tuning and malicious content generation. Therefore, we call for attention to these risks and encourage the adoption of defensive techniques to ensure the protection of personal content while advancing the development of generative AI technologies.
A.11 Limitations
@@
A.12 Multyview Camera Parameters
Building upon the dataset opensourced by Diffusion4D [13], Table 8 presents camera trajectory parameters, which serve as the foundation for consistent 4D content generation and subsequent reconstruction tasks.